# The influence of spin in black hole triplets

### v3

@@ 4 Astrophysical Implications
It is with no question that these binary systems are far less complicated than a triplet system. Tracing the orbits and predicting the spin based on frame-dragging effects are difficult but feasible for binaries. For the triplet system, these effects, especially on the more non-hierarchical triangles appear to be greatly chaotic and unpredictable. The overall effect of the rotating black hole appears negligible when considering the final states of these systems i.e. mergers and two-body encounters. But, utilizing the orbits of these bodies and the potentially observable LTE on the other two bodies of these types of systems to glean information about the spin magnitude of a black hole may be quite difficult. The system may not be constrained to the orbital plane and orbital elements oscillate largely, all of which contribute to unpredictable outcomes.
For triplets of SMBH, which are expected to form hierarchically via galactic mergers [13], this type of orbital element oscillation may not appear, which can limit unpredictability. This coincides with simulations as hierarchical triangles merge much more quickly, at smaller mass units, than the non-hierarchical triangles and therefore behave less chaotically. SBH triplets in the hierarchical formation however may experience this type of motion. Black holes of masses 1 M- 3 M may be highly theoretical but recently, the smallest black hole, of mass between 1.6 and 4.4 M been discovered via gravitational lensing [77, 76]. Raidal [61] also studied primordial black holes in the mass range of 0.1 M-103 M and their merger rates for potential LIGO detections. In some cases they found that a single primordial black hole nearby the binary would be the main disruption of a merger event. Here we find that for the most part, small mass triplet systems end in escape over merger with a remnant binary pair.
@@
Conclusion
We have studied the effect of spin in black hole triplets by conducting two sets of simulations. In Set One we looked at the Agekian-Anosova homology region (or region D) and in Set Two we looked at an ensemble of Pythagorean triangles. The results of Set One show that $\sim$22of simulations change how they end their lifetimes (from merger to escape or vice versa) with the inclusion of spin in this region. In Set Two, we have found that while there is not a significant effect of spin on some parameters like the fraction of mergers, the lifetimes of these systems may have been affected, particularly in the intermediary mass range (104 M-107 M). We have also found that certain effects are more visible in the non-hierarchical triangles than in the hierarchical triangles like the motion of the orbital elements $\omega$ and $\Omega$, owing to the chaotic tendencies of the non-hierarchical configurations. The inclusion of spin which is not perpendicular to the orbital plane leads to 3D motion.
@@ Appendix A Agekian-Anosova Region Extended
To fully cover all the possible geometrical configurations, we also conduct simulations with spinning black hole first, initialised at (0.5, 0), Run Three, and then initialised at (-0.5, 0), Run Four. In both cases, the masses of all the bodies remain as previously described and the spin magnitude on the spinning black hole also is kept at the same value. In the scenario where the spinning black hole is kept fixed on the positive axis, the log(lifetime/crossing time) of the systems appear Figure 13. Longer living systems cluster to the top left – away from the more massive rotating black hole, while systems that start off closer to it (bottom right) end very quickly.
Figure 13: The log(lifetime/crossing time) for simulations in the Agekian Anosova region with spin on M3 when it is initialised on the positive x-axis.
For the case where the spinning black hole is kept fixed on the negative axis, the log(lifetime/crossing time) looks like Figure 14. The results in this case appear much different than in the other cases as most of the simulations are longer living. This is most likely due to the positioning of the most massive, rotating hole further away from the other two black holes as compared to the previous cases.
Figure 14: The log(lifetime/crossing time) for simulations in the Agekian Anosova region with spin on M3 when it is initialised on the negative x-axis.

### v4

@@ 4 Discussion and Conclusion
It is with no question that these binary systems are far less complicated than a triplet system. Tracing the orbits and predicting the spin based on frame-dragging effects are difficult, but feasible for binaries. For the triplet system, these effects, especially on the more non-hierarchical triangles appear to be greatly chaotic and unpredictable. The overall effect of the rotating black hole appears negligible when considering the final states of these systems i.e. mergers and two-body encounters. But, utilizing the orbits of these bodies and the potentially observable LTE on the other two bodies of these types of systems to glean information about the spin magnitude of a black hole may be quite difficult. The system may not be constrained to the orbital plane and orbital elements oscillate largely, all of which contribute to unpredictable outcomes.
For triplets of SMBH, which are expected to form hierarchically via galactic mergers [13], this type of orbital element oscillation may not appear, which can limit unpredictability. This coincides with simulations as hierarchical triangles merge much more quickly, at smaller mass units, than the non-hierarchical triangles and therefore behave less chaotically. SBH triplets in the hierarchical formation however may experience this type of motion. Black holes of masses 1 M- 3 M are considered to be highly theoretical. There have been a few candidates for the smallest black hole discovered. One of these was found by Lam [76] via gravitational lensing, where the object could either be the smallest black hole found yet or a neutron star of mass 1.6 -- 4.4 M. When Sahu [77] observed this object however they estimated a mass of around 7 M. Lam [78] reanalysed this object and found that it is closer to 6 M. Raidal [61] also studied primordial black holes in the mass range of 0.1 M-103 M and their merger rates for potential LIGO detections. In some cases they found that a single primordial black hole nearby the binary would be the main disruption of a merger event. Here we find that for the most part, small mass triplet systems end in escape over merger with a remnant binary pair.
@@ Appendix A Agekian-Anosova Region Extended
To fully cover all the possible geometrical configurations, we also conduct simulations with spinning black hole first, initialised at (-0.5, 0), Run Three, and then initialised at (0.5, 0), Run Four. In both cases, the masses of all the bodies remain as previously described and the spin magnitude on the spinning black hole also is kept at the same value.
For the case where the spinning black hole is kept fixed on the negative axis, the log10(lifetime/crossing time) looks like Figure 12. The results in this case appear much different than in the other cases as most of the simulations are longer living. This is most likely due to the positioning of the most massive, rotating hole further away from the other two black holes as compared to the previous cases.
Figure 12: The log10(lifetime/crossing time) for simulations in the Agekian Anosova region for Run Three of Set One simulations. See Table 1.
In the scenario where the spinning black hole is kept fixed on the positive axis, the log10(lifetime/crossing time) of the systems appear Figure 13. Longer living systems cluster to the top left – away from the more massive rotating black hole, while systems that start off closer to it (bottom right) end very quickly.
Figure 13: The log10(lifetime/crossing time) for simulations in the Agekian Anosova region for Run Four of Set One simulations. See Table 1.